OpenStudy (anonymous):

a right angle has an acute angle θ such that sinθ=5/8. find the exact value of tanθ

OpenStudy (zzr0ck3r):

soh cah toa

OpenStudy (maheshmeghwal9):

\[\sin \theta = \frac{perpendicular}{hypotenuse}= \frac{5}{8}\]use pythagorus theorem & find base then \[\tan \theta = \frac{perpendicular}{base}= ??\]

OpenStudy (maheshmeghwal9):

\[h^2=p^2+b^2.\rightarrow Pythagorus \space theorem.\]

OpenStudy (mimi_x3):